# Hamiltonian simulation-based quantum-selected configuration interaction for large-scale electronic structure calculations with a quantum computer

### v2

@@ Abstract
Quantum-selected configuration interaction (QSCI) is an approach for quantum chemical calculations using current quantum computers. In conventional QSCI, Slater determinants used for the wave function expansion are sampled by iteratively performing approximate wave function preparation and subsequent measurement in the computational basis, and then the subspace Hamiltonian matrix is diagonalized on a classical computer. In this approach, preparation of a high-quality approximate wave function is necessary to accurately compute total energies. Here we propose a Hamiltonian simulation-based QSCI (HSB-QSCI) to avoid this difficulty, by sampling the Slater determinants from quantum states generated by the real-time evolution of approximate wave functions. We provide numerical simulations for the lowest spin-singlet and triplet states of oligoacenes (benzene, naphthalene, and anthracene), phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene. We found that the HSB-QSCI is applicable not only to molecules where the Hartree–Fock provides a good approximation of the ground state, but also to strongly correlated systems where preparing a high-quality approximate wave function is hard. Hardware demonstrations of the HSB-QSCI are also reported for carbyne molecules expressed by up to 36 qubits, using an IBM Quantum processor. The HSB-QSCI captures more than 99.75% of the correlation energies in the active space by considering about 1% of all the Slater determinants in 36 qubit systems, illustrating the ability of the proposed method to efficiently consider important electronic configurations.
@@ I Introduction
Among the diverse topics in the field of quantum computing, quantum chemical calculations of atoms and molecules have attracted attention as a promising application of quantum computers. A method to perform the full-configuration interaction (full-CI) calculation using the quantum phase estimation (QPE) algorithm was proposed in 2005 [1], and proof-of-principle experiments for the full-CI/STO-3G of the H2 molecule were reported in 2010 [2, 3]. The quantum circuit for QPE-based full-CI is usually very deep, so highly accurate QPE demonstrations were limited to small models with a few qubits [4, 5, 6, 7, 8]. Recently a QPE-type algorithm has been demonstrated in models with up to 33 qubits [9], however, it is still challenging to compute total energies of larger molecules with chemical precision (an error less than 1.0 kcal mol-1) on quantum computers available today.
@@
To reduce the computational load of quantum computers, several quantum–classical hybrid algorithms have been proposed. In 2014, the variational quantum eigensolver (VQE) was proposed for quantum chemical calculations using noisy intermediate-scale quantum (NISQ) [10] devices. In VQE, the quantum state corresponding to an approximate wave function is generated using a parameterized quantum circuit, and the energy expectation value is computed by repeatedly performing state preparation and measurements. The parameters are then optimized to minimize the energy expectation value on a classical computer. These steps are iterated until convergence is reached. VQE has been extensively studied both theoretically and experimentally [11], but it often suffers from challenges related to the high sampling cost required to achieve chemical precision [12] and issues with variational optimization in the presence of barren plateaus [13].
As an alternative approach to quantum chemical calculations using current quantum devices, the quantum-selected configuration interaction (QSCI) method has been proposed [14]. In the QSCI, quantum computers are used to sample Slater determinants that are important contributors to the ground state wave function. This can be done by running a quantum circuit to prepare an approximate wave function and measuring the quantum state in the computational basis. The subspace Hamiltonian matrix is then constructed from the Slater determinants selected via quantum computation and diagonalized on a classical computer. In this approach, the quantum computer is anticipated to efficiently sample a polynomial number of important Slater determinants from an exponentially large Hilbert space, and the nature of the quantum states used for sampling determines the accuracy of the calculation. At the stage of preparing an approximate wave function in QSCI, we may employ VQE or the adiabatic state preparation method [14]. To advance the former technique, combining an adaptive strategy for ansatz construction (ADAPT [15]) to enhance the efficiency of VQE [16], and using a simpler cost function for VQE parameter optimization [17] have also been proposed. However, these approaches may still face challenges in variational optimizations when implemented on quantum hardware. It should also be noted that in the latter approach [17], the VQE wave function with optimized parameters does not necessarily correspond to the minimum with respect to the given Hamiltonian.
@@
Another strategy to prepare the quantum state for QSCI is to perform the coupled cluster singles and doubles (CCSD) calculation on a classical computer and embed the CCSD wave function using the local unitary cluster Jastrow (LUCJ) ansatz [18]. By using the LUCJ ansatz and introducing an error mitigation technique called as a self-consistent configuration recovery (SCCR), which is inspired by the structure of chemistry problems, Robledo-Moreno and coworkers reported the QSCI calculations on the IBM superconducting quantum processor and supercomputer “Fugaku” for the nitrogen molecule (N2) and the iron–sulfur clusters [2Fe–2S] and [4Fe–4S] with up to 77 qubits [19]. This approach has also been applied to interaction energy [20] and excited state calculations [21, 22]. The combination of the QSCI with density matrix embedding theory has also been reported [23]. While this approach appears to be successful, it is not clear whether it is superior to the method of sampling important Slater determinants from the CCSD wave function on a classical computer.
@@
In this work, we report a Hamiltonian simulation-based quantum state preparation for QSCI (HSB-QSCI), in which important Slater determinants are sampled from the quantum states after the real-time evolutions of initial wave functions 111After posting this manuscript to arXiv, Mikkelsen and Nakagawa, and Yu and coworkers independently submitted their own manuscripts to arXiv, presenting the same idea explored in this study. M. Mikkelsen, Y. O. Nakagawa, arXiv:2412.13839, 2024; J. Yu, et al, arXiv:2501.09721, 2025.. Our proposed approach has several advantages over existing methods: (1) A simple approximate wave function, such as the HF wave function, can be used as the initial wave function, and no variational optimization in VQE is required. (2) Different samples can be obtained by changing the duration of the evolution time. (3) The accuracy of the Hamiltonian simulation does not need to be very high. As we will show later, the first-order Trotter decomposition with the time length of a single Trotter step is sufficient to collect important Slater determinants for the organic molecules under study. In addition, Hamiltonian term truncation based on operator locality in the localized molecular orbital basis can help to reduce the computational cost. (4) Feedback of the QSCI results into quantum computations is possible; that is, Slater determinants found to be important in the QSCI wave function can be added to the initial wave functions. (5) Although short-time evolution can be accurately simulated on a classical computer, long-time evolution poses significant challenges for classical simulation.
Note that methods of sampling basis states for Hamiltonian matrix diagonalization based on Hamiltonian simulation have already been investigated in the framework of quantum Krylov subspace algorithms [25, 26]. Unlike the quantum Krylov method, which computes and on a quantum computer by sampling, the HSB-QSCI method requires only measurements in the computational (Pauli-Z) basis and is thus less computationally demanding. Here, , where , is the initial wave function, and is the Hamiltonian of the system. In the context of the Monte Carlo simulation, a quantum dynamics Hamiltonian Monte Carlo (qdHMC) method is also proposed [27], in which the proposal step is performed on a quantum computer by sampling after time evolution.
@@
As a proof-of-concept demonstration of the HSB-QSCI method, we performed numerical simulations for the spin-singlet ground (S0) state and the first excited spin-triplet (T1) state of oligoacenes [benzene (1; 12 qubits), naphthalene (2; 20 qubits), and anthracene (3; 28 qubits)], phenylene-1,4-dinitrene (4; 20 qubits), and hexa-1,2,3,4,5-pentaene (5; 20 qubits) in both planar and twisted geometries. Our numerical simulations showed that the HSB-QSCI can calculate the total energy with chemical precision by sampling the Slater determinants with about 10–20 steps of Hamiltonian simulations. We also report hardware demonstrations of the HSB-QSCI for the S0 and T1 states of three carbyne (one-dimensional sp-hybridized carbon atom chain) molecules including 5, octa-1,2,3,4,5,6,7-heptaene (6; 28 qubits), and deca-1,2,3,4,5,6,7,8,9-nonaene (7; 36 qubits) using an IBM Quantum processor ibm_kawasaki, with the aid of matrix product operator (MPO)-based classical compression of the quantum circuit for time evolution. The HSB-QSCI energies calculated on ibm_kawasaki agree with the CAS-CI values within 0.6 kcal mol-1 of error in 5. In 6, chemical precision for the total energy was not achieved with the real quantum device, but the HSB-QSCI predicted the singlet–triplet energy gap and the energy difference between planar and twisted geometries in chemical precision. The energy difference between planar and twisted geometries in the S0 state of 7 was also calculated with an error of 0.29 kcal mol-1. In 7, the HSB-QSCI captures more than 99.75% of correlation energies by considering only about 1% of all the Slater determinants.
@@ II Theory
Here, is the HF determinant and are electronically excited determinants from the HF, and are the corresponding expansion coefficients. In the CI method, the number of Slater determinants included in the wave function expansion increases rapidly with the number of molecular orbitals and electrons, and the excitation order. To calculate the total energy with high accuracy and a low computational cost, methods to perform the CI expansion with selected Slater determinants have been investigated [28, 29, 30, 31]. The QSCI method uses a quantum computer to select important Slater determinants through the approximate wave function preparation and the subsequent computational-basis measurement. In the HSB-QSCI method, important Slater determinants are sampled from the measurement of the quantum state obtained from the time evolution of an approximate wave function.
@@ III Computational conditions
To demonstrate the HSB-QSCI, we performed numerical simulations on a classical computer. The target systems are the S0 and T1 states of oligoacenes (1, 2, and 3), 4 as a representative system of the open-shell singlet ground state, and 5, as shown in Figure 2. We also performed hardware demonstrations using an IBM Quantum processor for carbyne molecules 5, 6, and 7. Geometry optimizations of oligoacenes and carbynes in planar geometry were performed at the B3LYP/6-31G* level of theory. The twisted geometry of carbynes is generated by rotating one of the CH2 moieties 90 degrees from the equilibrium geometry. The geometry optimization of 4 was carried out at the CASSCF(10e,10o)/6-31G* level. Here, (e,o) represents the active space consisting of electrons and molecular orbitals. Cartesian coordinates of all molecules are given in Supporting Information.
@@
Since the computational cost of numerical simulations of quantum circuits scales exponentially with the number of qubits, we adopted the active space approximation. We used the (6e,6o), (10e,10o), and (14e,14o) active spaces for 1, 2, and 3, respectively, which consist of the valence and orbitals. The one- and two-electron integrals of oligoacenes are calculated at the RHF/STO-3G level. The active space of 4 contains the valence and orbitals and the in-plane 2p orbital of the nitrogen atoms, and the size is (10e,10o). The CASSCF(10e,10o)/6-31G* molecular orbitals are used to construct the second quantized Hamiltonian of 4. In the case of carbyne molecules 5, 6, and 7 with (10e,10o), (14e,14o), and (18e,18o) active spaces, respectively, the RHF/6-31G* calculation is performed and then the occupied molecular orbitals are localized using the Pipek–Mezey method [37]. For the virtual orbitals, we first formed the singular value decomposition (SVD) quasi-atomic external orbitals using an SVD with respect to the accurate atomic minimal basis functions, and then formed the ordered external orbitals using exchange integrals (using a keyword EXTLOC=ATMNOS in GAMESS-US software [38]). The active orbitals of all molecules are provided in Figures S1–S10 in Supporting Information. One- and two-electron atomic integrals are obtained from GAMESS-US, and the AO–MO integral transformations were performed using our own Python3 program. The qubit Hamiltonian is then generated using the OpenFermion library [39]. All the DFT calculations were done with the Gaussian 16 package [40], and the RHF and CASSCF calculations were carried out with the GAMESS-US software.
@@
To sample the Slater determinants based on the real-time evolution of the initial wave function, we set the evolution time for to be and , unless otherwise noted. To construct the quantum circuit for , we adopt the first-order Trotter decomposition with a single Trotter slice. Magnitude ordering [41] is used for the Trotterized term ordering. The quantum circuit simulations were done using the qsim library [42], which allows us to use GPGPU. The subspace Hamiltonian matrix construction and diagonalization was done using PyCI library [43], and the SCCR for hardware demonstrations was carried out using qiskit-addon-sqd library [44]. It should be noted that the subspace Hamiltonian diagonalization can also be done using qiskit-addon-sqd, but it uses a selected CI kernel_fixed_space subroutine in PySCF [45] where the spin and occupancies are stored separately, and the Slater determinants constructed from all possible combinations of spin and strings are used as the basis for the wave function expansion. In this implementation, Slater determinants not sampled by the quantum computer can be included in the subspace Hamiltonian diagonalization, resulting in a quadratic increase in the dimension of the Hamiltonian matrix. This can be seen as a kind of subspace enlargement at the expense of compactness of the QSCI wave function. We also performed a subspace Hamiltonian diagonalization using PySCF, and the results are given in Supporting Information section D. Quantum circuit simulations and the subspace Hamiltonian diagonalization in the QSCI part were performed on the NVIDIA DGX H100 system.

### v3

@@ Abstract
Quantum-selected configuration interaction (QSCI) is an approach for quantum chemical calculations using current quantum computers. In conventional QSCI, Slater determinants used for the wave function expansion are sampled by iteratively performing approximate wave function preparation and subsequent measurement in the computational basis, and then the subspace Hamiltonian matrix is diagonalized on a classical computer. In this approach, preparation of a high-quality approximate wave function is necessary to accurately compute total energies. Here we propose a Hamiltonian simulation-based QSCI (HSB-QSCI) to avoid this difficulty, by sampling the Slater determinants from quantum states generated by the real-time evolution of approximate wave functions. We provide numerical simulations for the lowest spin-singlet and triplet states of oligoacenes (benzene, naphthalene, and anthracene), phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene. We found that the HSB-QSCI is applicable not only to molecules where the Hartree–Fock provides a good approximation of the ground state, but also to strongly correlated systems where preparing a high-quality approximate wave function is hard. Hardware demonstrations of the HSB-QSCI are also reported for carbyne molecules expressed by up to 36 qubits, using an IBM Quantum processor. The HSB-QSCI captures more than 99.18% of the correlation energies in the active space by considering about 1% of all the Slater determinants in 36 qubit systems, illustrating the ability of the proposed method to efficiently consider important electronic configurations.
@@ I Introduction
Among the diverse topics in the field of quantum computing, quantum chemical calculations of atoms and molecules have attracted attention as a promising application of quantum computers. A method to perform the full-configuration interaction (full-CI) calculation using the quantum phase estimation (QPE) algorithm was proposed in 2005 [1], and proof-of-principle experiments for the full-CI/STO-3G of the H2 molecule were reported in 2010 [2, 3]. The quantum circuit for QPE-based full-CI is usually very deep, so highly accurate QPE demonstrations were limited to small models with a few qubits [4, 5, 6, 7, 8]. Recently a QPE-type algorithm has been demonstrated in models with up to 33 qubits [9], however, it is still challenging to compute total energies of larger molecules with chemical precision on quantum computers available today, where chemical precision is defined as an error from the full-CI energy [or complete active space configuration interaction (CAS-CI) energy when the active space approximation is employed] less than 1.0 kcal mol-1.
@@
To reduce the computational load of quantum computers, several quantum–classical hybrid algorithms have been proposed. In 2014, the variational quantum eigensolver (VQE) was proposed for quantum chemical calculations using noisy intermediate-scale quantum (NISQ) devices [10]. In VQE, the quantum state corresponding to an approximate wave function is generated using a parameterized quantum circuit, and the energy expectation value is computed by repeatedly performing state preparation and measurements. The parameters are then optimized to minimize the energy expectation value on a classical computer. These steps are iterated until convergence is reached. VQE has been extensively studied both theoretically and experimentally [11], but it often suffers from challenges related to the high sampling cost required to achieve chemical precision [12] and issues with variational optimization in the presence of barren plateaus [13].
As an alternative approach to quantum chemical calculations using current quantum devices, the quantum-selected configuration interaction (QSCI) method has been proposed [14]. In the QSCI, quantum computers are used to sample Slater determinants that are important contributors to the ground state wave function. This can be done by running a quantum circuit to prepare an approximate wave function and measuring the quantum state in the computational basis. The subspace Hamiltonian matrix is then constructed from the Slater determinants selected via quantum computation and diagonalized on a classical computer. Note that the concept of the selected CI has been well investigated in classical computation, and various selected CI methods have been developed, such as the CI using a perturbative selection made iteratively (CIPSI) [15], heat-bath CI (HCI) [16], and adaptive sampling CI (ASCI) [17]. These methods are often combined with multi-reference perturbation theory [18, 19]. In the quantum domain, QSCI deals with the variational optimization of the wave function in the selected subspace, and the effects of the remaining dynamical correlations can be taken into account by combining it with, for example, auxiliary-field quantum Monte Carlo (AFQMC) [20, 21] or general multi-configurational quasi-degenerate perturbation theory (GMC-QDPT) [22], although the perturbative energy correction from the QSCI is outside the scope of this study.
@@
In the QSCI approach, the quantum computer is anticipated to efficiently sample a polynomial number of important Slater determinants from an exponentially large Hilbert space, and the nature of the quantum states used for sampling determines the accuracy of the calculation. At the stage of preparing an approximate wave function in QSCI, we may employ VQE or the adiabatic state preparation method [14]. To advance the former technique, combining an adaptive strategy for ansatz construction (ADAPT [23]) to enhance the efficiency of VQE [24], and using a simpler cost function for VQE parameter optimization [25] have also been proposed. However, these approaches may still face challenges in variational optimizations when implemented on quantum hardware. It should also be noted that in the latter approach [25], the VQE wave function with optimized parameters does not necessarily correspond to the minimum with respect to the given Hamiltonian. Another strategy to prepare the quantum state for QSCI is to perform the coupled cluster singles and doubles (CCSD) calculation on a classical computer and embed the CCSD wave function using the local unitary cluster Jastrow (LUCJ) ansatz [26]. By using the LUCJ ansatz and introducing an error mitigation technique called as a self-consistent configuration recovery (SCCR), which is inspired by the structure of chemistry problems, Robledo-Moreno and coworkers reported the QSCI calculations on the IBM superconducting quantum processor and supercomputer “Fugaku” for the nitrogen molecule (N2) and the iron–sulfur clusters [2Fe–2S] and [4Fe–4S] with up to 77 qubits [27]. This approach has also been applied to interaction energy [28] and excited state calculations [29, 30]. The combination of the QSCI with density matrix embedding theory has also been reported [31]. While this approach appears to be successful, it is not clear whether it is superior to the method of sampling important Slater determinants from the CCSD wave function on a classical computer.
@@
In this work, we report a Hamiltonian simulation-based quantum state preparation for QSCI (HSB-QSCI), in which important Slater determinants are sampled from the quantum states after the real-time evolutions of initial wave functions [32]. Our proposed approach has several advantages over existing methods: (1) A simple approximate wave function, such as the HF wave function, can be used as the initial wave function, and no variational optimization in VQE is required. (2) Different samples can be obtained by changing the duration of the evolution time, and higher-order excitations can be considered even for short-time evolution. (3) The accuracy of the Hamiltonian simulation does not need to be very high. As we will show later, the first-order Trotter decomposition with the time length of a single Trotter step in atomic unit is sufficient to collect important Slater determinants for the organic molecules under study. In addition, Hamiltonian term truncation based on operator locality in the localized molecular orbital basis can help to reduce the computational cost. (4) Feedback of the QSCI results into quantum computations is possible; that is, Slater determinants found to be important in the QSCI wave function can be added to the initial wave functions. (5) Although short-time evolution can be accurately simulated on a classical computer, long-time evolution poses significant challenges for classical simulation.
Note that methods of sampling basis states for Hamiltonian matrix diagonalization based on Hamiltonian simulation have already been investigated in the framework of quantum Krylov subspace (QKS) algorithms [33, 34]. In the QKS method, the Hamiltonian matrix and the overlap matrix are computed using a quantum computer via a Hadamard test, and the generalized eigenvalue problem is solved on a classical computer. Here, , where and is the initial wave function. Since the Taylor expansion of the time evolution operator yields polynomials of as follows, time evolution is expected to help finding electron configurations that are important for the wave function expansion.
Unlike HSB-QSCI, QKS is applicable to systems in which the correlated wave function is supported by an exponentially large number of Slater determinants. However, the overlap matrix becomes ill-conditioned when the subspace basis is nearly linearly dependent [34, 35], and the sampling cost in QKS becomes large in order to avoid numerical instability. In contrast, HSB-QSCI requires only measurements in the computational (Pauli-Z) basis and therefore the computational cost on a quantum computer is significantly reduced. HSB-QSCI is also expected to be numerically more stable, because Slater determinants are orthonormal, while subspace Hamiltonian diagonalization on a classical computer can be a bottleneck. In the context of the Monte Carlo simulation, a quantum dynamics Hamiltonian Monte Carlo (qdHMC) method is also proposed [36], in which the proposal step is performed on a quantum computer by sampling after time evolution.
@@
As a proof-of-concept demonstration of the HSB-QSCI method, we first carried out numerical simulations for the spin-singlet ground state of the H2O molecule, and compared the results obtained using different sampling strategies (randomly sampling Slater determinants preserving spatial symmetry, , and , where and represent the spin magnetic quantum number and the number of electrons, respectively, VQE-based QSCI [14], and HCI). Then, we performed numerical simulations for the spin-singlet ground (S0) state and the first excited spin-triplet (T1) state of oligoacenes [benzene (1; 12 qubits), naphthalene (2; 20 qubits), and anthracene (3; 28 qubits)], phenylene-1,4-dinitrene (4; 20 qubits), and hexa-1,2,3,4,5-pentaene (5; 20 qubits) in both planar and twisted geometries. Our numerical simulations showed that the HSB-QSCI can calculate the total energy with chemical precision by sampling the Slater determinants with about 10–20 steps of Hamiltonian simulations. We also report hardware demonstrations of the HSB-QSCI for the S0 and T1 states of three carbyne (one-dimensional sp-hybridized carbon atom chain) molecules including 5, octa-1,2,3,4,5,6,7-heptaene (6; 28 qubits), and deca-1,2,3,4,5,6,7,8,9-nonaene (7; 36 qubits) using an IBM Quantum processor ibm_kawasaki, with the aid of matrix product operator (MPO)-based classical compression of the quantum circuit for time evolution. The HSB-QSCI energies calculated on ibm_kawasaki agree with the CAS-CI values within 0.6 kcal mol-1 of error in 5. In 6, chemical precision for the total energy was not achieved with the real quantum device, but the HSB-QSCI predicted the singlet–triplet energy gap and the energy difference between planar and twisted geometries in chemical precision. The energy difference between planar and twisted geometries in the S0 state of 7 was also calculated with an error of 0.29 kcal mol-1. In 7, the HSB-QSCI captures more than 99.18% of correlation energies by considering only about 1% of all the Slater determinants.
@@ II Theory
Here, is the HF determinant and are electronically excited determinants from the HF, and are the corresponding expansion coefficients. In the CI method, the number of Slater determinants included in the wave function expansion increases rapidly with the number of molecular orbitals and electrons, and the excitation order. To calculate the total energy with high accuracy and a low computational cost, methods to perform the CI expansion with selected Slater determinants have been investigated [15, 16, 37, 38, 17]. The QSCI method uses a quantum computer to select important Slater determinants through the approximate wave function preparation and the subsequent computational-basis measurement. In the HSB-QSCI method, important Slater determinants are sampled from the measurement of the quantum state obtained from the time evolution of an approximate wave function.
@@ III Computational conditions
To demonstrate the HSB-QSCI, we performed numerical simulations on a classical computer. The target systems are the S0 state of H2O molecule, the S0 and T1 states of oligoacenes (1, 2, and 3), 4 as a representative system of the open-shell singlet ground state, and 5, as shown in Figure 2. We also performed hardware demonstrations using an IBM Quantum processor for carbyne molecules 5, 6, and 7. For the H2O molecule, we used the same geometry as in the original QSCI paper [14]. Geometry optimizations of oligoacenes and carbynes in planar geometry were performed at the B3LYP[44, 45]/6-31G*[46, 47] level of theory. In all DFT calculations, we used default settings in Gaussian 16 [48]. The SCF convergence thresholds were set to and for the RMS and maximum density changes, respectively, and Hartree for the energy changes. For geometry optimizations, the convergence criteria were 0.000450 and 0.000300 for the maximum and RMS forces, and 0.001800 and 0.001200 for the maximum and RMS displacements, respectively. We used UltraFine grid with a pruned (99,590 grid), having 99 radial shells and 590 angular points per shell. The twisted geometry of carbynes is generated by rotating one of the CH2 moieties 90 degrees from the equilibrium geometry. The geometry optimization of 4 was carried out at the CASSCF(10e,10o)/6-31G* level. Here, (e,o) represents the active space consisting of electrons and molecular orbitals. For the CASSCF geometry optimization, we used a gradient convergence threshold of Hartree/Bohr, which is the default setting in the GAMESS-US [49] software. Convergence is achieved when the largest gradient component is below the specified threshold and the root mean square gradient is less than 1/3 of the threshold. Cartesian coordinates of all molecules are given in Supporting Information.
@@
Since the computational cost of numerical simulations of quantum circuits scales exponentially with the number of qubits, we adopted the active space approximation. For the H2O molecule, we used the (6e,5o) active space, consistent with the QSCI study by Kanno and coworker [14]. We used the (6e,6o), (10e,10o), and (14e,14o) active spaces for 1, 2, and 3, respectively, which consist of the valence and orbitals. The one- and two-electron integrals of H2O and oligoacenes are calculated at the RHF/STO-3G [50] level. The active space of 4 contains the valence and orbitals and the in-plane 2p orbital of the nitrogen atoms, and the size is (10e,10o). The CASSCF(10e,10o)/6-31G* molecular orbitals are used to construct the second quantized Hamiltonian of 4. In the case of carbyne molecules 5, 6, and 7 with (10e,10o), (14e,14o), and (18e,18o) active spaces, respectively, the RHF/6-31G* calculation is performed and then the occupied molecular orbitals are localized using the Pipek–Mezey method [51]. For the virtual orbitals, we first formed the singular value decomposition (SVD) quasi-atomic external orbitals using an SVD with respect to the accurate atomic minimal basis functions, and then formed the ordered external orbitals using exchange integrals (using a keyword EXTLOC=ATMNOS in GAMESS-US software [49]). The active orbitals of all molecules are provided in Figures S1–S10 in Supporting Information. One- and two-electron atomic orbital integrals are obtained from GAMESS-US, and the AO–MO integral transformations were performed using our own Python3 program. The qubit Hamiltonian is then generated using the OpenFermion library [52]. All the DFT calculations were done with the Gaussian 16 package [48], and the RHF and CASSCF calculations were carried out with the GAMESS-US software [49].
@@
To sample the Slater determinants based on the real-time evolution of the initial wave function, we set the evolution time for to be in atomic unit and , unless otherwise noted. To construct the quantum circuit for , we adopt the first-order Trotter decomposition with a single Trotter slice. Magnitude ordering [53] is used for the Trotterized term ordering. The quantum circuit simulations were done using the qsim library [54], which allows us to use GPGPU. The subspace Hamiltonian matrix construction and diagonalization was done using PyCI library [55], and the SCCR for hardware demonstrations was carried out using qiskit-addon-sqd library [56]. It should be noted that the subspace Hamiltonian diagonalization can also be done using qiskit-addon-sqd, but it uses a selected CI kernel_fixed_space subroutine in PySCF [57] where the spin and occupancies are stored separately, and the Slater determinants constructed from all possible combinations of spin and strings are used as the basis for the wave function expansion. In this implementation, Slater determinants not sampled by the quantum computer can be included in the subspace Hamiltonian diagonalization, resulting in a quadratic increase in the dimension of the Hamiltonian matrix. This can be seen as a kind of subspace enlargement at the expense of compactness of the QSCI wave function. We also performed a subspace Hamiltonian diagonalization using PySCF, and the results are given in Supporting Information section D. Quantum circuit simulations and the subspace Hamiltonian diagonalization in the QSCI part were performed on the NVIDIA DGX H100 system.